5. 1.5 mLs of 6 M HCl are added to 250 mL of water (pH=7.0). What is the resulting (H) and the final pH?
Answer #1

MIVM2 V2 = MV 6x15 M2x50 -Mx51.5 ph-Log LH -PH M2- CH= 16 M 6x1.5+ 167x250 25-5 9 +o.000025 255 M = o-035 7/ [Hl Total00 357